a man whose mass is 60kg rides a ferris wheel that has a radius of 20m. The man has a scale available in the cart he is riding. If hi speed during the ride is a constant 5m/s, what would the scale read at the top of the ride?

(a) 588 N

(b) 513 N

(c) 663 N

(d) 680 N

(e) 703 N

please explain!
Answer #1

Gravitational acceleration = g = 9.8 m/s2

Mass of the man = m = 60 kg

Radius of the Ferris wheel = R = 20 m

Speed of the ride = V = 5 m/s

The reading on the scale is the normal force exerted by the seat on the man.

Normal force on the man = N

When the man is at the top of the ride there are only two forces acting on the man, his own weight downwards and the normal force upwards.

For the circular motion of the man there needs to be a centripetal force acting downwards when the man is the at the top of the ride that is the net force due to the weight of the man and the normal force on him is equal to the centripetal force.

mV2/R = mg - N

(60)(5)2/(20) = (60)(9.8) - N

N = 513 N

Reading of the scale at the top of the ride = 513 N